OpenStudy (anonymous):

hE lP !! !! Tell whether the sequence is arithmetic. If it is, what is the common difference? 15, 19, 23, 27 . . . . A) yes ; 15/19 B) yes ; 4 C) yes ; 34 D) no

OpenStudy (alexandervonhumboldt2):

in an arafmetic secuence difference beetwen terms is the same

OpenStudy (alexandervonhumboldt2):

19-15=23-18=27-23

OpenStudy (anonymous):

B Difference is 4

OpenStudy (anonymous):

B. 4 is difference.. :)

OpenStudy (anonymous):

so B people ?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

thankya :D